OpenStudy (anonymous):

find all normal subgroups of {Z6, +}

OpenStudy (anonymous):

any subgroup of \[\{\mathbb Z_6,+\}\] will be normal because it is abelian.

OpenStudy (anonymous):

normal subgroups will be the trivial subgroup \[\{0\}\] the subgroup of order 2 isomorphic to \[\mathbb Z_2\] \[\{0,3\}\] and the subgroup of order 3 isomorphic to \[\mathbb Z_3\] \[\{0,2,4\}\] and the whole group

OpenStudy (anonymous):

alright, good to know, abelian groups always have normal subgroups. thanks

OpenStudy (anonymous):

that is any subgroup of an abelian group is normal. i am not sure what definition of normal you are using, but if it is \[x^{-1}hx\in H \forall \text{ }h\in H \] then if it is abelian you see that \[x^{-1}hx=x^{-1}xh=eh=h\] so all subgroups must be normal